Science Report – April 27th

Science Report for 27 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No locations were scouted during the EVA today. Emergency triage procedures were practiced so that a crew member might be safely transported to an evacuation location in the event of an incident.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar flare in the direction of our planet and immediately alerted mission control to the danger.

3. Shortwave Texting on EVA’s: Hand signal communication was used during the emergency EVA practice today. Further practice is needed to achieve a smooth process.

4. GPS Route Measurements: No measurements today, as the EVA was at the hab location.

5. Waterless dish cleaning: Samples were taken from a dirty dish and a dirty dish that had been put under the UV light. There are minimal signs of bacterial growth from the first set of samples. It may take a couple more days to see signs of growth. Both sets of samples are now in the incubator in the hab.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Analysis will be done as the mission continues.

Copyright © The Mars Society. All rights reserved. | Main Site